OpenStudy (anonymous):

Help Please! I have no clue how to do this. What is the graph of the function. f(x)=1/3x^2

OpenStudy (anonymous):

Oh these. Eh...you could always just plug in a set of values for X and then solve for f(x) to get the multiple coordinates. Then just connect them with a line.

OpenStudy (anonymous):

So what do you mean im confused. sorry. @JoyLin

OpenStudy (anonymous):

Algebra isn't really my strong point.

OpenStudy (anonymous):

Here I'll do the first step with you. Let x=1. Then plug it in. F(x)=(1/3)(1)^2 F(x)=(1/3)(1) F(x)=1/3 Your coordinates are now: X=1, f(x)=1/3 (1, 1/3)

OpenStudy (anonymous):

It's ok! Don't worry! Math can get a little confusing at times.

OpenStudy (anonymous):

Now let's try 2. X=2 F(x)=(1/3)(2)^2 F(x)=(1/3)(4) F(xx)=4/3 The second coordinates are X=2 F(x)=4/3 (2, 4/3)

OpenStudy (anonymous):

So graph (1, 1/3) and (2, 4/3) as points.

OpenStudy (anonymous):

Im sure you can guess what's next. Make x=3

OpenStudy (anonymous):

See if you can do it. I just want to make sure you understand it.

OpenStudy (anonymous):

oh!!!! cX thank you so much!

OpenStudy (anonymous):

You get the general idea. :) Hopefully that helped! Remember, learning is asking and messing up, so don't be afraid to do either! Ever need my help again? tag or message. Until then, good luck!